# TE-PAI: Exact Time Evolution by Sampling Random Circuits

### v2

@@ II.2 Gate count in the random circuits
We explicitly bound the leading constant factors in the error term in $\mathbb{E}(\nu)=\nu_{\infty}+O(N^{-1})$ in our proof in Appendix C. We thus find that for a constant rotation angle $\Delta$, TE-PAI saturates the Lieb-Robinson bound [32, 33], which determines a fundamental bound on the speed at which local information can spread due to time evolution under local interactions. However, we will later prove that in practice, $\Delta$ needs to scale with both $T$ and the system size to avoid an exponential increase in the sample complexity.
@@ II.3 Measurement overhead
In TE-PAI, we randomly replace the continuous-angle gates with three discrete gate variants: the third gate variant has a very low associated probability for small $\Delta$, however, when it does get selected then any measured observable is multiplied with a negative sign. This negative sign leads to an increase in the variance of the expectation value of the observable being estimated. Thus, in order to estimate the observable expected value to the same precision as with an infinitely deep Trotter circuit, one needs to perform an increased number of measurements. We detail in Appendix B that this measurement overhead is upper bounded by the following factor as
@@
We explicitly derive and bound the leading constant factors in the above error term in our proof in Appendix D. We note that a variant of PAI was developed in ref [39] that introduces a trade-off parameter $\lambda$ that allows to continuously interpolate between the unbiased PAI estimator (exact estimator with measurement overhead $\|g\|_{1}$) and an approximate, biased estimator which has no measurement overhead at all. The numerical approach of ref [39] can be used immediately for reducing the above measurement overhead at the cost of introducing a bias, however, for ease of notation in the present work we focus on the exact, unbiased implementation.
We illustrate in Fig. 3, that the measurement overhead of our unbiased estimator grows exponentially with the total time $T$ but decreasing $\Delta$ results in a slower exponential blowup. Furthermore, in Fig. 4, we illustrate the trade-off between the measurement overhead and the expected number of gates at different rotation angles $\Delta$. While the expected number of gates $\nu_{\infty}$ is a constant that is independent of $N$, decreasing the rotation angle $\Delta$ increases $\nu_{\infty}$ and ultimately can yield to divergence in the limit $\lim_{\Delta\rightarrow 0}\mathbb{E}(\nu)=\infty$.
@@
Time complexity: The end-to-end time complexity of estimating an expected value to precision $\epsilon$ in TE-PAI scales according to standard shot-noise scaling as $O(T^{2}\epsilon^{-2}Q^{-1}\exp[Q])$. We can compare this to a first-order Trotter circuit from which the expected value is extracted using amplitude estimation in which case the Trotter circuit is repeated coherently $O(\epsilon^{-1})$ times leading to a total time complexity $O(T^{2}\epsilon^{-2})$.
@@
In conclusion, the time complexity of extracting a time-evolved expected value in TE-PAI is comparable to using first-order Trotterisation in combination with amplitude estimation. However, TE-PAI has a number of significant advantages. First, TE-PAI requires only shallow circuits, making it feasible to run when coherence time or code distance is limited, whereas amplitude amplification requires significantly deeper circuits. Second, while amplitude estimation estimates observables one-by-one, we detail below that TE-PAI is compatible with advanced measurement techniques, including classical shadows and thus allows simultaneous estimation of many observables. Third, TE-PAI can be fully parallelized, with the sampling task distributed across many quantum computers, whereas some amplitude estimation variants can be parallelised, but only to a more limited extent [40]. Finally, TE-PAI only requires the implementation of a single type of non-Clifford rotation, specifically a single-qubit rotation with angle $\Delta$, which allows us to design a particularly efficient fault-tolerant implementation below.
@@
Figure 6: (left) Distribution of estimated expected values $\langle X_{0}\rangle$ from time evolution circuits that implement evolution under the 14-qubit spin Hamiltonian in Eq. 13 for $T=1$ and using $N_{s}=1000$ circuit repetitions (shots). (left, grey) distribution of $\langle X_{0}\rangle$ from a deep Trotter circuit with $N=1000$ layers, consisting of $\nu=56000$ continuous-angle rotations. (left, red) distribution of $\langle X_{0}\rangle$ from a shallow Trotter circuit consisting of $N=50$ layers and $\nu=2800$ continuous-angle rotations. (left, blue) Histogram and estimated distribution of $\langle X_{0}\rangle$ using TE-PAI using only two kinds of discrete rotation angles $\Delta=2^{-7}\pi$ and $\pi$. The distribution width of TE-PAI is slightly increased but its mean is identical to that of the deep Trotter circuit (blue) while its gate count is very close to that of the shallow Trotter circuit as $\nu_{\infty}=2715$. (right) Expectation values $\langle X_{0}\rangle$ as the simulation time $T$ increases from 00 to $2$. Due to its shallow depth, the $N=100$ Trotter circuit (right, red) introduces a bias to the expected value measurement, albeit its standard deviation remains unchanged as we increase $T$. In contrast, TE-PAI (right, blue) recovers the same mean value as the $N=2000$ deep Trotter circuit (right, grey), however, its standard deviation increases as $T$ increases.
@@ III Numerical demonstrations
where we choose time-dependent coupling terms $J(t)=\cos(99\pi t)$ and the parameters $\omega_{k}$ are chosen uniformly randomly within the range $[-1,1]$. This model is representative of problems considered in condensed matter physics for studying many-body localisation. These problems could be effectively explored using early quantum computers and may be hard to simulate classically for large numbers of qubits [41, 42]. For the present demonstration, we use $n=14$ qubits and due to the periodicity of the Hamiltonian, $\overline{\|c\|_{1}}\approx 33.30$ is constant for integer evolution times $T=1,2\dotsc$, and we choose $T=1$ and $\Delta=2^{-7}\pi$. We generate random circuits using TE-PAI for a product formula of $N=1000$ layers and a total evolution time of $1$ – we chose a relatively low value of $N$ as in our demonstrations we will use a relatively low number of shots $N_{s}=1000$ and therefore shot noise will dominate over residual algorithmic errors (while indeed $N$ can be increased without requiring more quantum resources).
@@
Second, we execute the TE-PAI circuits to estimate expected values $\langle X_{0}\rangle$ and compare them to conventional Lie-Trotter circuits. In Fig. 6 (left) we report the distribution of expected values estimated using $N_{s}=1000$ shots: while TE-PAI in Fig. 6 (left, blue) has a slightly increased distribution width, its mean value matches exactly the mean value of an arbitrarily deep Trotter circuit in Fig. 6 (left, grey). Furthermore, TE-PAI only uses a small number of gates comparable to a shallow Trotter circuit ($N=50$) – which shallow circuit introduces a significant bias due to significant algorithmic errors, as shown in Fig. 6 (left, red).
@@
Finally, we generate a family of TE-PAI circuits for an increasing total simulation time $T$ and compare the expected values $\langle X_{0}\rangle(T)$ to shallow and deep Trotter circuits in Fig. 6 (right). Expected values estimated from shallow Trotter circuits have relatively low statistical uncertainty Fig. 6 (right, red error bars) throughout the evolution, however, suffer from a significant bias, i.e., red curve is significantly off from the reference deep time evolution circuit Fig. 6 (right, grey). In contrast, TE-PAI (blue line) achieves the same mean value as the deep Trotter circuit, however, its statistical uncertainty increases with the simulation time.
@@
III.1 Noisy Implementation
In Fig. 7, we repeat simulations of the spin-ring model defined in Section III but assuming a noisy 7-qubit system and a rotation angle $\Delta=\frac{\pi}{2^{6}}$ which yields an expected number of gates of $1364$ at $T=2$. We consider Trotterisation using $N=100$ (green) and $N=200$ layers (magenta) and observe that increasing the number of Trotter layers introduces a more significant bias due to the increased number of noisy gates. In contrast, TE-PAI achieves a smaller bias, i.e., blue is closer to the reference simulation (grey), as it requires fewer noisy quantum gates.
Additionally, to further reduce the effect of gate noise, TE-PAI can naturally be combined with quantum error mitigation [43] and can also be combined with classical shadows as detailed in ref. [16]. Furthermore, TE-PAI estimates an expectation value by executing a large number of structurally radically different random circuits; such randomisation protocols have been shown to scramble local gate noise to global depolarising noise—with theoretical proofs for global random circuits [44] and numerical evidence for shallow structured circuits [45]—which allows for very simple and effective error mitigation by global rescaling.
@@ IV Fault-tolerant resource estimation
TE-PAI implements time evolution exactly by averaging outputs of random circuits. A significant advantage compared to, e.g., Trotterisation, is that our circuits only have two kinds of rotation angles as ±$\pm\Delta$ and $\pi$, and, as we demonstrate in the following, this significantly reduces overall non-Clifford resources required for fault-tolerant implementations.
In particular, early generations of fault-tolerant quantum computers will likely be limited by the number of logical qubits and by the achievable circuit depths. Here, we perform resource estimation for a typical example of simulating the time evolution under a 100-qubit spin Hamiltonian (which we introduced in Section III). This Hamiltonian has $L=400$ terms, and we fix $T=1$, $\overline{\|c\|_{1}}=241.3$. The expected number of rotation gates using $\Delta=\frac{\pi}{2^{8}}$ is approximately $39,328$ and the measurement overhead is then about $19.32$.
In Appendix F, we detail how our random circuits can be compiled into a sequence of Clifford gates and discrete angle, single-qubit $Z$ rotation gates $R_{Z}(\Delta)$. Thus the only non-Clifford resource we require are single-qubit $Z$ rotation gates, all with the same rotation angle. In fault-tolerant quantum computing (FTQC), Clifford gates are relatively cheaper and less error prone than non-Clifford resources. Thus, our focus in this section is to minimize the implementation cost of our method by efficiently implementing the $R_{Z}(\Delta)$ rotations.

### v3

@@ II.2 Gate count in the random circuits
We explicitly bound the leading constant factors in the error term in $\mathbb{E}(\nu)=\nu_{\infty}+O(N^{-1})$ in our proof in LABEL:app:proof2. We thus find that for a constant rotation angle $\Delta$, TE-PAI saturates the Lieb-Robinson bound [LR, Haah2021], which determines a fundamental bound on the speed at which local information can spread due to time evolution under local interactions. However, we will later prove that in practice, $\Delta$ needs to scale with both $T$ and the system size to avoid an exponential increase in the sample complexity.
@@ II.3 Measurement overhead
In TE-PAI, we randomly replace the continuous-angle gates with three discrete gate variants: the third gate variant has a very low associated probability for small $\Delta$, however, when it does get selected then any measured observable is multiplied with a negative sign. This negative sign leads to an increase in the variance of the expectation value of the observable being estimated. Thus, in order to estimate the observable expected value to the same precision as with an infinitely deep Trotter circuit, one needs to perform an increased number of measurements. We detail in LABEL:app:PAI that this measurement overhead is upper bounded by the following factor as
@@
We explicitly derive and bound the leading constant factors in the above error term in our proof in LABEL:app:proof1. We note that a variant of PAI was developed in ref [koczor2024sparse] that introduces a trade-off parameter $\lambda$ that allows to continuously interpolate between the unbiased PAI estimator (exact estimator with measurement overhead $\|g\|_{1}$) and an approximate, biased estimator which has no measurement overhead at all. The numerical approach of ref [koczor2024sparse] can be used immediately for reducing the above measurement overhead at the cost of introducing a bias, however, for ease of notation in the present work we focus on the exact, unbiased implementation.
We illustrate in Fig. 3, that the measurement overhead of our unbiased estimator grows exponentially with the total time $T$ but decreasing $\Delta$ results in a slower exponential blowup. Furthermore, in Fig. 4, we illustrate the trade-off between the measurement overhead and the expected number of gates at different rotation angles $\Delta$. As we increase $\Delta$, we decrease the circuit depth but also increase the measurement overhead. While the expected number of gates $\nu_{\infty}$ is a constant that is independent of $N$, decreasing the rotation angle $\Delta$ increases $\nu_{\infty}$ and ultimately can yield to divergence in the limit $\lim_{\Delta\rightarrow 0}\mathbb{E}(\nu)=\infty$. In a quantum computer where the quantum gates are not perfect—for example, noisy physical operations or logical operations in an early fault-tolerant machine—the circuit depth needs to be chosen such that $p_{\mathrm{err}}^{-1}\propto\nu$ to ensure that error-mitigation techniques can be used effectively. This desired circuit depth can be achieved by setting the corresponding $\Delta$ angle.
Both the measurement overhead and the circuit depth of TE-PAI depends implicitly on the number of qubits, through the dependence of the norm $\overline{\lVert c\rVert}_{1}$ on the number of qubits $n$. For example, for the Heisenberg spin model considered later in this manuscript, $\overline{\lVert c\rVert}_{1}\in O(n)$, while in quantum chemistry a pessimistic worst-case bound is $\overline{\lVert c\rVert}_{1}\in O(n^{4})$. However, advanced techniques such as tensor hypercontraction (THC) can significantly reduce this scaling [hyper_contraction]. The complexity of other quantum simulation methods, such as qubitization and qDRIFT, similarly depends implicitly on the number of qubits via the norm $\overline{\lVert c\rVert}_{1}$.
@@
Time complexity: We now estimate the end-to-end time complexity of TE-PAI. The number of gates in a single circuit is $O(T^{2}Q^{-1})$, and estimating an observable to precision $\epsilon$ requires $O(\epsilon^{-2}\exp{Q})$ repetitions (we assume the circuits are repeated serially and therefore TE-PAI does not require a space overhead). The end-to-end time complexity of TE-PAI (or equivalently its space-time volume) therefore scales as $O(T^{2}\epsilon^{-2}Q^{-1}\exp[Q])$. We can compare this to a first-order Trotter circuit from which the expected value is extracted using amplitude estimation in which case the Trotter circuit is repeated coherently $O(\epsilon^{-1})$ times leading to a total time complexity $O(T^{2}\epsilon^{-2})$ (the space-time volume of this approach scales similarly given amplitude estimation requires little space overhead).
@@
In conclusion, the time complexity of extracting a time-evolved expected value in TE-PAI is comparable to using first-order Trotterisation in combination with amplitude estimation. However, TE-PAI has a number of significant advantages. First, TE-PAI requires only shallow circuits, making it feasible to run when coherence time or code distance is limited, whereas amplitude amplification requires significantly deeper circuits. Second, while amplitude estimation estimates observables one-by-one, we detail below that TE-PAI is compatible with advanced measurement techniques, including classical shadows and thus allows simultaneous estimation of many observables. Third, certain platforms, such as silicon qubit technologies, enable the fabrication of many independent QPUs on a single chip with no or limited communication between QPUs. TE-PAI can be fully parallelized, with the sampling task distributed across many QPUs, leading to a proportional reduction of the runtime. In contrast, some amplitude estimation variants can be parallelised, but only to a more limited extent [labib2024quantum]. Finally, TE-PAI only requires the implementation of a single type of non-Clifford rotation, specifically a single-qubit rotation with angle $\Delta$, which allows us to design a particularly efficient fault-tolerant implementation below.
Note that, in terms of gate complexity, truncated Taylor–series LCU methods scale as $\mathcal{O}\!\left(\|c\|_{1}\,T\,\frac{\log(1/\epsilon)}{\log\log(1/\epsilon)}\right)$ [LCU1, LCU2, LCU3], while qubitization, QSP, and QSVT scale as $\mathcal{O}\!\left(\|c\|_{1}\,T+\log(1/\epsilon)\right)$ [QSP1, QSP2, QSVT], but these techniques are fundamentally different and require deep coherent circuits that are not feasible to execute on current or early-FTQC devices unlike TE-PAI.
@@
Figure 6: (left) Distribution of estimated expected values $\langle X_{0}\rangle$ from time evolution circuits that implement evolution under the 14-qubit spin Hamiltonian in Eq. 13 for $T=1$ and using $N_{s}=1000$ circuit repetitions (shots). (left, grey) distribution of $\langle X_{0}\rangle$ from a deep Trotter circuit with $N=1000$ layers, consisting of $\nu=56000$ continuous-angle rotations. (left, red) distribution of $\langle X_{0}\rangle$ from a shallow Trotter circuit consisting of $N=50$ layers and $\nu=2800$ continuous-angle rotations. (left, blue) Histogram and estimated distribution of $\langle X_{0}\rangle$ using TE-PAI using only two kinds of discrete rotation angles $\Delta=2^{-7}\pi$ and $\pi$. The distribution width of TE-PAI is slightly increased but its mean is identical to that of the deep Trotter circuit (blue) while its gate count is very close to that of the shallow Trotter circuit as $\nu_{\infty}=2715$. (right) Expectation values $\langle X_{0}\rangle$ as the simulation time $T$ increases from 0 to $2$. Due to its shallow depth, the $N=100$ Trotter circuit (right, red) introduces a bias to the expected value measurement, albeit its standard deviation remains unchanged as we increase $T$. In contrast, TE-PAI (right, blue) recovers the same mean value as the $N=2000$ deep Trotter circuit (right, grey), however, its standard deviation increases as $T$ increases.
@@ III.1 Heisenberg spin model benchmark
where we choose time-dependent coupling terms $J(t)=\cos(99\pi t)$ and the parameters $\omega_{k}$ are chosen uniformly randomly within the range [1,1][-1,1]. This model is representative of problems considered in condensed matter physics for studying many-body localisation. These problems could be effectively explored using early quantum computers and may be hard to simulate classically for large numbers of qubits [PhysRevB.91.081103, Childs_2018]. For the present demonstration, we use $n=14$ qubits and due to the periodicity of the Hamiltonian, $\overline{\|c\|_{1}}\approx 33.30$ is constant for integer evolution times $T=1,2\dotsc$, and we choose $T=1$ and $\Delta=2^{-7}\pi$. We generate random circuits using TE-PAI for a product formula of $N=1000$ layers and a total evolution time of $1$ – we chose a relatively low value of $N$ as in our demonstrations we will use a relatively low number of shots $N_{s}=1000$ and therefore shot noise will dominate over residual algorithmic errors (while indeed $N$ can be increased without requiring more quantum resources).
@@
Second, we execute the TE-PAI circuits to estimate expected values $\langle X_{0}\rangle$ and compare them to conventional Lie-Trotter circuits. In Fig. 6 (left) we report the distribution of expected values estimated using $N_{s}=1000$ shots: while TE-PAI in Fig. 6 (left, blue) has a slightly increased distribution width, its mean value matches exactly the mean value of an arbitrarily deep Trotter circuit in Fig. 6 (left, grey). Furthermore, with small measurement overhead of $\|g\|_{1}^{\infty}\approx 2.15$, TE-PAI requires only a small number of gates comparable to a shallow Trotter circuit with $N=50$ – which shallow circuit introduces a significant bias due to significant algorithmic errors, as shown in Fig. 6 (left, red).
@@
Finally, we generate a family of TE-PAI circuits for an increasing total simulation time $T$ and compare the expected values $\langle X_{0}\rangle(T)$ to shallow and deep Trotter circuits in Fig. 6 (right). Since this plot considers $T=0\ldots 2$, we doubled the Trotter step numbers to $N=2000$ for the deep Trotter circuit and $N=100$ for the shallow Trotter circuit. Expected values estimated from shallow Trotter circuits have relatively low statistical uncertainty Fig. 6 (right, red error bars) throughout the evolution, however, suffer from a significant bias, i.e., red curve is significantly off from the reference deep time evolution circuit Fig. 6 (right, grey). In contrast, TE-PAI (blue line) achieves the same mean value as the deep Trotter circuit, however, its statistical uncertainty increases with the simulation time.
III.2 Quantum chemistry benchmark
The complexity of TE-PAI depends on the coefficient $\ell_{1}$-norm $\overline{\|c\|_{1}}$ rather than on the number of Hamiltonian terms $L$. This suggests that TE-PAI could be advantageous for quantum chemistry problems with large $L$, where standard Trotter methods require prohibitively large gate counts per Trotter step. An important example is the simulation of the FeMoCo active-space Hamiltonian. TE-PAI is therefore expected to play a role in enabling quantum advantage in challenging quantum chemistry applications.
As a simple demonstration, we consider a time-independent chemistry simulation of a linear chain of $N_{\mathrm{H}}=6$ hydrogen atoms in the STO-6G minimal basis with an interatomic spacing of $2.0$ Bohr, described by the corresponding electronic Hamiltonian:
$H=\sum_{pq}h_{pq}\,a^{\dagger}_{p}a_{q}+\frac{1}{2}\sum_{pqrs}h_{pqrs}\,a^{\dagger}_{p}a^{\dagger}_{q}a_{r}a_{s},$ (14)
where $a_{p}^{\dagger}$ and $a_{q}$ are fermionic creation and annihilation operators, and $h_{pq}$, $h_{pqrs}$ denote one- and two-electron integrals obtained from a restricted Hartree-Fock calculation. Mapping to qubits is performed using the Jordan-Wigner transformation, which yields a Pauli expansion acting on twelve qubits (corresponding to twelve spin orbitals).
Since orbital occupation numbers,
$n_{p}=a^{\dagger}_{p}a_{p}=\tfrac{1-Z_{p}}{2},$
are an important observable in quantum chemistry, we compute the single-qubit expectation value $\langle Z_{0}(t)\rangle$, which is directly related to the occupation of the first spin orbital.
[…]
In chemistry simulations typically a relatively high precision, the chemical accuracy $\propto 10^{-3}$ is required which is challenging to meet using shallow Trotter circuits. TE-PAI, on the other hand, reproduces the exact result, with the only remaining error coming from sampling noise, which can be reduced by increasing the number of shots.
[…]
III.3 Noisy Implementation
In Fig. 8, we repeat simulations of the spin-ring model defined in Section III but assuming a noisy 7-qubit system and a rotation angle $\Delta=\frac{\pi}{2^{6}}$ which yields an expected number of gates of $1364$ at $T=2$. We consider Trotterisation using $N=100$ (green) and $N=200$ layers (magenta). For both $N=100$ and $N=200$, the Trotter step sizes are larger than those of the shallow Trotter circuit in Fig. 6, but they are still small enough that the Trotter error cannot be ignored in our noiseless simulations. Despite the small number of steps, the resulting circuits remain highly susceptible to noise. We also observe that increasing the number of Trotter layers introduces a more significant bias due to the increased number of noisy gates. In contrast, TE-PAI achieves a smaller bias, i.e., blue is closer to the reference simulation (grey), as it requires fewer noisy quantum gates. This configuration of $\Delta$ enables us to flexibly reduce the circuit depth, thereby enhancing its resilience to noise compared with Trotter circuits.
Additionally, to further reduce the effect of gate noise, TE-PAI can naturally be combined with quantum error mitigation [RevModPhys.95.045005] and can also be combined with classical shadows as detailed in ref. [mitigation_shadow]. Furthermore, TE-PAI estimates an expectation value by executing a large number of structurally radically different random circuits; such randomisation protocols have been shown to scramble local gate noise to global depolarising noise—with theoretical proofs for global random circuits [dalzell2024random] and numerical evidence for shallow structured circuits [white_noise]—which allows for very simple and effective error mitigation by global rescaling.
@@ IV Fault-tolerant resource estimation
TE-PAI implements time evolution exactly by averaging outputs of random circuits. A significant advantage compared to, e.g., Trotterisation, is that our circuits only have two kinds of rotation angles as $\pm\Delta$ and $\pi$, and, as we demonstrate in the following, this significantly reduces overall non-Clifford resources required for fault-tolerant implementations.
In particular, early generations of fault-tolerant quantum computers will likely be limited by the number of logical qubits and by the achievable circuit depths. Here, we perform resource estimation for a typical example of simulating the time evolution under a 100-qubit spin Hamiltonian (which we introduced in Section III). This Hamiltonian has $L=400$ terms, and we fix $T=1$, $\overline{\|c\|_{1}}=241.3$. The expected number of rotation gates using $\Delta=\frac{\pi}{2^{8}}$ is approximately $39,328$ and the measurement overhead is then about $19.32$. We set the Trotter step to $N=10^{4}$ because, by Eq. 3 in LABEL:app:classical_cost, this choice guarantees a Trotter error to be $\epsilon_{T}\leq 0.1$. However, such worst-case bounds are typically loose by about an order of magnitude [trotter_err], so we expect a error on the order of $10^{-2}$ in practice.
In LABEL:app:resource, we detail how our random circuits can be compiled into a sequence of Clifford gates and discrete angle, single-qubit $Z$ rotation gates $R_{Z}(\Delta)$. Thus the only non-Clifford resource we require are single-qubit $Z$ rotation gates, all with the same rotation angle. In fault-tolerant quantum computing (FTQC), Clifford gates are relatively cheaper and less error prone than non-Clifford resources. Thus, our focus in this section is to minimize the implementation cost of our method by efficiently implementing the $R_{Z}(\Delta)$ rotations.
Fault-tolerant machines are emerging; however, there is a range of different hardware platforms with radically different qubit technologies. For example, solid-state platforms (superconducting qubits and quantum dots) have fast gates but relatively noisier qubits, and the leading error-correction approach is the surface code. On the other hand, ion traps and neutral atoms have significantly higher fidelities but slower gate times, and therefore the requirements for error correction can be quite different. Nevertheless, most platforms aim to implement fault-tolerant Clifford gates augmented with magic state teleportation, and therefore we expect the results in this section to apply to a broad range of hardware platforms.